Niagara Falls are 50n high. Calculate the potential energy of 0.1cubic meter of water at the top relative to the bottom. Density of water is 1000kgm^-3. g=10m/s^2
1
Expert's answer
2019-04-18T09:57:44-0400

The potential energy

"PE=mgh=\rho V g h"

"=1000\:\rm{kg /m^3}\times 0.1\;\rm{m^3}\times 10\;\rm{m/s^2}\times 50\;\rm{m}=50 000 \:\rm{J}""=50\:\rm{kJ}"